OpenStudy (anonymous):

Simplify the expression √18 + √50 A. 3√2 B. 5√2 C 8√2 D. 34√2 Please explain how to do this, don't just give the answer.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

explain please...

OpenStudy (anonymous):

\[\sqrt{25}\times \sqrt{2}+\sqrt{9}\times \sqrt{2}\]

OpenStudy (anonymous):

First find the factors for the numbers inside the radical For 18, we have 2 x 3 x 3 U do the other

OpenStudy (anonymous):

\[5\sqrt{2}+3\sqrt{2}\]

OpenStudy (anonymous):

2 x 5 x 5

OpenStudy (anonymous):

ty levinoyo